The difference between 100 and me is 86. What numbers can I be using one and two step problems involving addition and subtraction
Answer by ikleyn(52814)

.
Two answer are possible and are equally valid: 186 and 14.


186 = 100 + 86,   and   14 = 100-86.


I don't know what it means "using one and two step problems involving addition and subtraction",

but I have a common sense in my mind, instead. And it is more than enough.